November 8-14, 1997
November 8 , 9 , 10 , 11 , 12 , 13
, 14
November 8
  - User run.
  
 - Operation status in this morning:
  
|   Time | 
  Beam current | 
  Beam lifetime | 
  I*tau | 
  Average pressure | 
  Integrated current | 
|   10:00 AM | 
  377.0 mA | 
  623 minutes | 
  235 A*minutes | 
  0.345 nTorr | 
  235.9 A*hours | 
   - We found a leak of insulation oil from one of the klystron
  power supplies. We made temporary repairs on it.
 

Beam current (blue) and beam lifetime (red)
around the clock.
November 9
  - User run.
  
 - Operation status in this morning:
  
|   Time | 
  Beam current | 
  Beam lifetime | 
  I*tau | 
  Average pressure | 
  Integrated current | 
|   10:00 AM | 
  376.3 mA | 
  635 minutes | 
  239 A*minutes | 
  0.329 nTorr | 
  242.9 A*hours | 
 

Beam current (blue) and beam lifetime (red)
around the clock.
November 10
  - Machine study.
  
 - We made the first trial operation of one of the low-emittance optics (phase
  advance: 90 degrees/cell). We succeeded
  to store the beam at 9:49 PM. 
  The new PF ring can be operated under three low-emittance optics, together
  with a medium emittance optics (currently used for user runs). The beam
  optics, used today, is the moderate one which corresponds to the design
  beam emittance of 44 nm*radians. Also see "High-brilliance
  project of the PF storage ring".
 

Image from the beam profile monitor which
showed successfull beam storege.
Beam current: 5.65 mA. 
  - The maximum beam current of 5.65 mA was stored under the low-emittance
  optics (90 degrees/cell) at 0:18 AM on Nov. 11.
 

Plot of the beam current (a signal from the
DC current transformer as shown by an yellow trace) showing the successful
beam storege under the low-emittance optics. Ordinate: beam current (10
mA/full), abscissa: time.
  - New orbit feedback system was successfully tested during
  3:50 - 8:26 AM on Nov. 11. (We fixed some problems that remained, and carried
  out final test). 
  The orbit feedback system has come
  into operation under user runs from Nov. 11 !
   - We also repaired the oil leak from a klystron power supply
  which occured on Nov. 8. (We thank Denkenseiki Corp. for their quick correspondence).
 

Beam current (blue trace) and beam lifetime
(red trace) around the clock.
November 11
  - User run.
  
 - Operation status in this morning: 
|   Time | 
  Beam current | 
  Beam lifetime | 
  I*tau | 
  Average pressure | 
  Integrated current | 
|   10:00 AM | 
  344.1 mA | 
  661 minutes | 
  227 A*minutes | 
  0.316 nTorr | 
  251.4 A*hours | 
   - The new orbit feedback
  system, which is used to correct vertical beam movements,
  has come into operation today.
 

Beam current (blue) and beam lifetime (red)
around the clock.
November 12
  - User run.
  
 - During the beam injection in this morning, we had a trouble
  with one of the power supplies for septum magnets (i.e. deflection magnets
  for injected beams). It took about three hours to look at and to repair
  it. The trouble was caused by a bad electrical contact around a knife switch
  for input AC line. We finished beam injection at 0:45 PM.
 

Beam current (blue trace) and lifetime (red
trace) around the clock.
November 13
  - User run.
  
 - A water leak was found in a plant building (so called
  "Energy Center") for the linac. Repair works took place today.
  Beam injection, scheduled in the evening, was moved to 9:00 PM due to this
  work. (The repair works were actually finished at 3:47 PM).
  
 - We had a beam dump at 9:22 PM, just after the beam injection.
  The beam was injected again.
  
 - Operation status in this morning: 
|   Time | 
  Beam current | 
  Beam lifetime | 
  I*tau | 
  Average pressure | 
  Integrated current | 
|   10:00 AM | 
  374.2 mA | 
  643 minutes | 
  241 A*minutes | 
  0.381 nTorr | 
  264.7 A*hours | 
 

Beam current (blue trace) and beam lifetime
(red one) around the clock.
November 14
  - User run.
  
 - Operation status in this morning: 
|   Time | 
  Beam current | 
  Beam lifetime | 
  I*tau | 
  Average pressure | 
  Integrated current | 
|   10:00 AM | 
  376.4 mA | 
  691 minutes | 
  260 A*minutes | 
  0.274 nTorr | 
  270.9 A*hours | 
   - We had a beam dump at 5:02 PM due to an interlock action
  (so-called "Channel Dump") with the beamline BL-10. This trouble
  was due to a fault with a solenoid valve of main beam shutter, and we replaced
  this part with a spare.
 

Beam current (blue trace) and beam lifetime
(red one) around the clock.

Proress in the vacuum conditioning due to
irradiation of SR.
Products of the beam current and the beam lifetime are shown by red dots.
Average vacuum pressures per unit stored current are shown by blue dots.